OpenStudy (hannahwaller5):

Will Give Medal! Jules kicks a soccer ball off the ground and into the air with an initial velocity of 25 feet per second. Assume the starting height of the ball is 0 feet. Approximately what maximum height does the soccer ball reach? 0.8 ft 1.6 ft 9.8 ft 19.6 ft

OpenStudy (august899):

Assuming that Jules kicks the ball completely vertically, then you can use one of the famous kinematics equations (I'll spare you the derivation here): (v-final)^2 - (v-initial)^2 = 2*a*height In your case, the ball's final velocity (at its maximum height) is zero! Your acceleration is -g = -32 feet per second per second, and your initial velocity is 25 feet per second. So you have: height = - (25)^2 / (2*-32) = 625 / 64 = 9.8 feet

OpenStudy (hannahwaller5):

Ok so if i solve - (25)^2 / (2*-32) = 625 / 64 = 9.8 feet i will get my answer...?

OpenStudy (august899):

Right,when you do 25^2 you get 625 then when you do 2 times -32 you get 64,divide that and you get 9.8.

OpenStudy (hannahwaller5):

Oh ok that makes since thanks so much! (:

OpenStudy (august899):

You're welcome! :)